OpenStudy (anonymous):

find the slope of the line that passes through the pair of points . (1,7) , (10,1)

OpenStudy (anonymous):

(x1,y1)=(1,7) (x2,y2)=(10,1) slope=y2-y1/x2-x1 slope=1-7/10-1 slope=-6/9 reduce slope=-2/3 lots of more solved problem like this in math911 www.math911.com

OpenStudy (agreene):

\[\Delta y / \Delta x\]